# Emergent Open-Vocabulary Semantic Segmentation from Off-the-shelf Vision-Language Models

### v2

@@ Abstract
From image-text pairs, large-scale vision-language models (VLMs) learn to implicitly associate image regions with words, which prove effective for tasks like visual question answering. However, leveraging the learned association for open-vocabulary semantic segmentation remains a challenge. In this paper, we propose a simple, yet extremely effective, training-free technique, Plug-and-Play Open-Vocabulary Semantic Segmentation (PnP-OVSS) for this task. PnP-OVSS leverages a VLM with direct text-to-image cross-attention and an image-text matching loss. To balance between over-segmentation and under-segmentation, we introduce Salience Dropout; by iteratively dropping patches that the model is most attentive to, we are able to better resolve the entire extent of the segmentation mask. PnP-OVSS does not require any neural network training and performs hyperparameter tuning without the need for any segmentation annotations, even for a validation set. PnP-OVSS demonstrates substantial improvements over comparable baselines (+29.4% mIoU on Pascal VOC, +13.2% mIoU on Pascal Context, +14.0% mIoU on MS COCO, +2.4% mIoU on COCO Stuff) and even outperforms most baselines that conduct additional network training on top of pretrained VLMs. Our codebase is at https://github.com/letitiabanana/PnP-OVSS.
Figure 1: Qualitative Results of PnP-OVSS + BLIP. Images are from Pascal Context and COCO stuff. The right columns and bottom rows show the ground-truth (GT); the rest are our results. Note accurate results even on complex (trees) and small objects (last column).
@@ 1 Introduction
Figure 2: Segmentation masks for elephant and train using (a) off-the-shelf cross-attention, (b) cross-attention + GradCAM, and (c) cross-attention + GradCam + Salience DropOut (§3.2). The naive cross-attention masks are too inclusive whereas GradCAM is too exclusive.
@@
In summary, we propose Plug-and-Play Open-vocabulary Semantic Segmentation (PnP-OVSS), an extremely simple and training-free framework to extract semantic segmentations from VLMs. At zero extra training cost, PnP-OVSS can be used with any pretrained VLM with text-to-image cross attention layer and an image-text matching loss. It has zero reliance on pixel-level annotations, including a validation set for hyperparameter tuning. At the same time, PnP-OVSS delivers excellent performance. It not only beats the training-free baseline with remarkable margins (+29.4% mIoU on Pascal VOC, +13.2% on Pascal Context, +14.0% on MS COCO, +2.4% on COCO Stuff), but also outperforms most recent techniques within the past two years that require extensive finetuning on top of the VLM pretraining.
@@ 3 Method
PnP-OVSS has four major steps. First, we extract a cross-attention salience map per object class from a VLM. Second, we sharpen the salience map by weighing it with the ITM gradient in the style of GradCAM. Third, we apply Salience DropOut that iteratively completes the salience maps. Fourth, we apply Dense CRF [29] for fine-grained adjustment. The process is illustrated in Fig. 3 and Fig. 4.
In the first step, we feed an image and a text prompt to the pretrained VLM and extract cross-attention maps. The text prompt is “A picture of [class 1] [class 2][class K]”, which includes all $K$ class names of interest in the dataset. The image is divided into $P\times P$ patches. The text prompt and the image first go through the modality-specific encoders respectively, followed by a cross-attention fusion module. In the cross-attention layers, the text encodings serve as query vectors and the image patche encodings are as key and value vectors.
@@
With the correct layer and attention head, we observe that this attention map, when normalized by softmax along the first dimension $K$, can provide passable semantic segmentations. However, it tends to include many patches unrelated to the class being segmented, leading to over-segmentation. Hence, we introduce two refinement steps, GradCAM and Salience DropOut, explained in §3.1 and §3.2. After these steps, we acquire aggregate salience maps for every class. Finally, we conduct local polish to the resultant salience maps using Dense CRF, as described in §3.3.
@@ 3.2 Salience DropOut
Since our task is zero-shot and open-vocabulary, we do not have prior knowledge of the classes present in the image. Hence, we sum up the salience maps $\widetilde{M}^{(k)}$ over all classes $k$, yielding a class-agnostic salience map $U^{(t)}$ for the $t^{\text{th}}$ Salience DropOut iteration, $U^{(t)}=\sum^{K}_{k=1}\widetilde{M}^{(k,t)},$ where $\widetilde{M}^{(k,t)}$ is the GradCAM salience map for class $k$ after the $t^{\text{th}}$ iteration. Next, we zero out the 50% of image patches with the highest values in $U^{(t)}$. On the remaining image, we compute the GradCAM salience maps $\widetilde{M}^{(k,t+1)}$ and their sum, $U^{(t+1)}$. Any image patch previously zeroed out will always receive zero salience in later iterations.
@@
We stop at four rounds of dropout, as almost all (93.75%) patches are removed beyond that point. The final output for each class $k$ is the sum over all salience maps across the four dropout iterations, $\hat{M}^{(k)}=\sum_{t=1}^{4}\widetilde{M}^{(k,t)}$.
@@ 4.1 Datasets and Implementation Details
[…]
BridgeTower $\times$ $\times$ 322 36.4 32.3 24.2 27.6 18.6
BridgeTower $\times$ $\times$ 336 35.3 32.4 24.2 27.6 18.0
BridgeTower $\times$ $\times$ 770 35.2 32.4 24.5 28.0 19.0
Table 3: Zero-shot semantic segmentation performance in mIoU. Group 3 contains the most similar baselines that serve as fair comparisons to PnP-OVSS. Groups 1 and 2 benefit from additional training, extra image-text data, and hyperparameter tuning on dense labels. We use the word “by” followed by a paper citation to indicate results of the same technique reported by different papers.  CLIPpy tests on 133 categories of COCO Stuff while we test all 171 classes of COCO Stuff. ViL-Seg is tested on subset of classes on the three datasets, as detailed in the supplementary.
@@
+ GradCam 21.6 17.5
       + Drop 1 25.1 19.8
       + Drop 2 26.5 20.6
       + Drop 3 27.0 20.9
       + Drop 4 27.2 20.9
       + Drop 4 + Blur 36.8 28.6
       + Drop 4 + Dense CRF 35.3 31.8
       + Drop 4 + Blur + Dense CRF 42.8 32.5
@@ 4.2 Main Results
PnP-OVSS exhibits excellent performance. Comparing with MaskClip [90] and Reco [58], the two methods that require no additional training and ground truth for hyperparameter tuning, on an equal-resolution basis, we attain +29.4% mIoU on Pascal VOC, +13.2% mIoU on Pascal Context, +14.0% mIoU on COCO Object, and +11.4% mIoU on ADE-20K. Further, PnP-OVSS surpasses all baselines in Group 2. On an equal-resolution basis, we achieve +13.7% mIoU on Pascal Voc, +9.6% on Pascal Context, +9.5% on COCO Object, +11.6% on COCO Stuff, and +12.8% on ADE-20K.
@@
When applied to BridgeTower [79], PnP-OVSS still surpasses all methods in Group 3 by 10.2% on Pascal VOC, 5.1% on Pascal Context, 8.5% on COCO Object, 0.4% on COCO Stuff, and 6.8% on ADE-20K. This showcases the plug-and-play ability of PnP-OVSS, which excels with different base networks.
@@ 4.4 Hyperparameter Sensitivity
We make the following observations. First, different layers and heads have drastic performance differences. The best-to-worst difference among all layers is 20, and that among heads in Layer 8 is 23.8, underscoring the importance of hyperparameter tuning. Second, the random search using the proposed reward function correctly identifies the best layer-head combination, even though Layer 8 is not the best layer based on average head performance. This indicates the effectiveness of our method.
@@ 5 Conclusions
We propose PnP-OVSS, which extracts the ability of semantic segmentation from opaque VLMs. PnP-OVSS is simple to use, requires no extra finetuning, and delivers high performance, exceeding not only all baselines that requires no finetuning, but also all baselines that do not use image-text pairs in finetuning. Its success hints at a new direction for open-vocabulary segmentation tasks leveraging large VLMs.
@@ 7 Additional Qualitative Results
We provide additional qualitative results of images from Pascal Context and COCO Stuff in Fig 6. We note that, in the third image of the third row, PnP-OVSS correctly recognizes four loaves of bread whereas the ground truth only annotates one.
The rightmost column of the bottom three rows, separated from the rest by a red dash line, are examples of failure cases. The example on top contains multiple small instances of the same class (people and surfboard), which are understandably hard. The middle example contains multiple instances of people, which causes difficulties for PnP-OVSS to cover all objects. In addition, images with a clutter of different objects and complex texture (as in the last image) often cause a drop in performance.
Figure 6: Qualitative Results of PnP-OVSS + BLIP. Images are from Pascal Context and COCO stuff. The bottom rows show the ground-truth (GT); the rest are our results. The last column of the last three rows, after the red dash line, shows failure cases.

### v3

@@ Abstract
From image-text pairs, large-scale vision-language models (VLMs) learn to implicitly associate image regions with words, which prove effective for tasks like visual question answering. However, leveraging the learned association for open-vocabulary semantic segmentation remains a challenge. In this paper, we propose a simple, yet extremely effective, training-free technique, Plug-and-Play Open-Vocabulary Semantic Segmentation (PnP-OVSS) for this task. PnP-OVSS leverages a VLM with direct text-to-image cross-attention and an image-text matching loss. To balance between over-segmentation and under-segmentation, we introduce Salience Dropout; by iteratively dropping patches that the model is most attentive to, we are able to better resolve the entire extent of the segmentation mask. PnP-OVSS does not require any neural network training and performs hyperparameter tuning without the need for any segmentation annotations, even for a validation set. PnP-OVSS demonstrates substantial improvements over comparable baselines (+26.2% mIoU on Pascal VOC, +20.5% mIoU on MS COCO, +3.1% mIoU on COCO Stuff and +3.0% mIoU on ADE20K). Our codebase is at https://github.com/letitiabanana/PnP-OVSS.
Figure 1: Qualitative Results of PnP-OVSS + BLIP. Images are from Pascal VOC and COCO Object. The right columns and bottom rows show the ground-truth (GT); the rest are our results. Note the good results on small objects like the frisbee and the tennis racket.
@@ 1 Introduction
Figure 2: Segmentation masks for elephant and giraffe using (a) off-the-shelf cross-attention, (b) cross-attention + GradCAM, and (c) cross-attention + GradCam + Salience DropOut (§3.2). The naive cross-attention masks are too inclusive whereas GradCAM is too exclusive.
@@
In summary, we propose Plug-and-Play Open-vocabulary Semantic Segmentation (PnP-OVSS), an extremely simple and training-free framework to extract semantic segmentations from VLMs. At zero extra training cost, PnP-OVSS can be used with any pretrained VLM with text-to-image cross attention layer and an image-text matching loss. It has zero reliance on pixel-level annotations, including a validation set for hyperparameter tuning. At the same time, PnP-OVSS delivers excellent performance. It not only beats the training-free baseline with remarkable margins (+26.2% mIoU on Pascal VOC, +20.5% mIoU on MS COCO, +3.1% mIoU on COCO Stuff and +3.0% mIoU on ADE20K), but also outperforms most recent techniques within the past two years that require extensive finetuning on top of the VLM pretraining.
@@ 3 Method
PnP-OVSS has five major steps. First, we prompt GPT4o to output the categories appear in each image. Second, we extract a cross-attention salience map for each predicted category from a VLM. Third, we sharpen the salience map by weighing it with the ITM gradient in the style of GradCAM. Fourth, we apply Salience DropOut that iteratively completes the salience maps. Finally, we apply Dense CRF [29] for fine-grained adjustment. The process is illustrated in Fig. 3 and Fig. 4.
Directly inputting all the labels from the dataset into the VLM could introduce unnecessary noise into the segmentation process. In the first step, we prompt GPT4o to filter the possible categories within each image. The prompt for GPT4o is: “Help me find as much and accurate as possible, categories appear in the image. The the available categories are {CATEGORIES}. Double check if there is anything missing. You may output unsure categories and give them lower probability. Output strictly in the format [id1: classname1, id2: classname2, …], [probability of class1 in percentage, probability of class2 in percentage, …] without other words. ” {CATEGORIES} being a placeholder for all the categories in a dataset. The categories output by GPT4o with higher than 70% probability are treated as the predicted classes for the image. In the second step, we feed an image and a text prompt to the pretrained VLM and extract cross-attention maps. The text prompt is “A picture of [class 1] [class 2][class K]”, which includes the $K$ GPT4o predicted categories of the image. The image is divided into $P\times P$ patches. The text prompt and the image first go through the modality-specific encoders respectively, followed by a cross-attention fusion module. In the cross-attention layers, the text encodings serve as query vectors and the image patche encodings are as key and value vectors.
@@
With the correct layer and attention head, we observe that this attention map can provide passable semantic segmentations. However, it tends to include many patches unrelated to the class being segmented, leading to over-segmentation. Hence, we introduce two refinement steps, GradCAM and Salience DropOut, explained in §3.1 and §3.2. After these steps, we acquire aggregate salience maps for every class. Finally, we conduct local polish to the resultant salience maps using Dense CRF, as described in §3.3.
@@ 3.2 Salience DropOut
We sum up the salience maps $\widetilde{M}^{(k)}$ over the predicted classes $K$, yielding a class-agnostic salience map $U^{(t)}$ for the $t^{\text{th}}$ Salience DropOut iteration, $U^{(t)}=\sum^{K}_{k=1}\widetilde{M}^{(k,t)},$ where $\widetilde{M}^{(k,t)}$ is the GradCAM salience map for class $k$ after the $t^{\text{th}}$ iteration.
Next, we zero out 10 image patches with the highest values in $U^{(t)}$. On the remaining image, we compute the GradCAM salience maps $\widetilde{M}^{(k,t+1)}$ and their sum, $U^{(t+1)}$. Any image patch previously zeroed out will always receive zero salience in later iterations.
@@
We stop at the third rounds of dropout. The final output for each class $k$ is the sum over all salience maps across the three dropout iterations, $\hat{M}^{(k)}=\sum_{t=1}^{4}\widetilde{M}^{(k,t)}$.
@@ 4.1 Datasets and Implementation Details
BridgeTower $\times$ $\times$ 336 42.4 25.3 30.4 15.7 14.8
Table 3: Zero-shot semantic segmentation performance in mIoU. Group 3 contains the most similar baselines that serve as fair comparisons to PnP-OVSS. Groups 1 and 2 benefit from additional training, extra image-text data, and hyperparameter tuning on dense labels. We use the word “by” followed by a paper citation to indicate results of the same technique reported by different papers.  CLIPpy tests on 133 categories of COCO Stuff and Reco tests on 27 super categories while we test all 171 classes of COCO Stuff. ViL-Seg is tested on subset of classes on the three datasets, as detailed in the supplementary.
@@
+ GradCam 20.2 14.4
       + Drop 1 21.6 14.9
       + Drop 2 22.4 15.2
       + Drop 3 22.9 15.3
       + Drop 3 + Blur 26.1 17.1
       + Drop 3 + Dense CRF 25.1 16.0
       + Drop 3 + Blur + Dense CRF 28.9 17.9
@@ 4.2 Main Results
PnP-OVSS exhibits excellent performance. Comparing with MaskClip [90] and Reco [58], the two methods that require no additional training and ground truth for hyperparameter tuning, on an equal-resolution basis, we attain +26.2% mIoU on Pascal VOC, +0.8% mIoU on Pascal Context, +20.5% mIoU on COCO Object, +3.1% mIoU COCO Stuff and +3.0% mIoU ADE-20K. Further, PnP-OVSS surpasses all baselines except MaskClip with ST in Group 2.
@@
When applied to BridgeTower [79], PnP-OVSS still surpasses all methods in Group 3 by 17.3% on Pascal VOC, 14.7% on COCO Object, 0.9% on COCO Stuff, and 3.6% on ADE-20K. This showcases the plug-and-play ability of PnP-OVSS, which excels with different base networks.
@@ 4.4 Hyperparameter Sensitivity
We make the following observations. First, different layers and heads have drastic performance differences. The best-to-worst difference among all layers is 24.4, and that among heads in Layer 8 is 30.3, underscoring the importance of hyperparameter tuning. Second, the random search using the proposed reward function correctly identifies the best layer-head combination This indicates the effectiveness of our method.
@@ 5 Conclusions
We propose PnP-OVSS, which extracts the ability of semantic segmentation from opaque VLMs. PnP-OVSS is simple to use, requires no extra finetuning, and delivers high performance, exceeding not only all baselines that requires no finetuning, but also many of the baselines that require finetuning. Its success hints at a new direction for open-vocabulary segmentation tasks leveraging large VLMs.
@@ 7 Additional Qualitative Results
We provide additional qualitative results of images from Pascal Context and COCO Stuff in Fig 6. We note that, in the third image of the third row, PnP-OVSS correctly recognizes the occluded person missing in the ground truth.
The bottom row are examples of failure cases. The examples on top contains multiple small instances of the same class (people and forks, which are understandably hard. All the example contains multiple instances of people, which causes difficulties for PnP-OVSS to cover all objects. In addition, images with a clutter of different objects and complex texture often cause a drop in performance.
Figure 6: Qualitative Results of PnP-OVSS + BLIP. Images are from Pascal Context, COCO Object and COCO stuff. The bottom rows show the ground-truth (GT); the rest are our results. The last row shows failure cases.